OpenStudy (anonymous):

For the following system, if you isolated y in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 2 x − y = −6 Choose one answer. A. −6 − x B. −6 + x C. 6 − x D. 6 + x

OpenStudy (anonymous):

In the second equation, first add y+6 to both sides. Then you will get y value in terms of x. Then substitute this y value in 1st equation to get your x

OpenStudy (anonymous):

x - y = -6 x = -6 + y x+6 = y D 6 + x

OpenStudy (anonymous):

@suzukirider1400 actually, you don't need to do the second step. First step will suffice for the given question . (in the steps I mentioned)